# Multi-Scale Cell Decomposition for Path Planning using Restrictive Routing Potential Fields

### v1

@@ V Conclusion
In this work, we have presented Larp, an innovative path planning framework tailored for restrictive routing within potential fields. Demonstrating superior results, Larp outperforms both traditional and contemporary potential field-based methodologies by reliably crafting safer routes that diminish the risk of restriction violations, all while preserving comparative efficiency in terms of travel distance. This is achieved through an initial segmentation of the potential field into multi-scale cells, with each cell’s restriction zone determined by its proximity to nearby obstacles. For transparency and to facilitate further research, the source code for an official implementation of Larp has been made publicly accessible.
While Larp marks an advancement in route safety, opportunities for improvement remain. The framework’s quad cell decomposition approach and zone classification system are poised for further optimization. The influence of a cell’s dimensions and positioning on the placement of subsequent cells is a critical factor in the pursuit of optimally short and safe routes. Moreover, the current method of zone classification, which relies on an proxy scaled distance to a cell’s periphery to gauge the highest potential across the region, may estimate an higher potential for a region that what it is. Future endeavors will concentrate on refining these elements and developing a dynamic and scalable system capable of storing and updating cells in real-time to adapt to evolving restrictions for unmanned aircraft system traffic management (UTM).

### v3

@@ V Conclusion
This study introduces Larp, a novel path planning framework specifically designed for restrictive routing within potential fields. Larp has demonstrated superior performance compared to both traditional and contemporary potential field-based methodologies by consistently devising safer routes that mitigate the risk of restriction violations while maintaining competitive efficiency in terms of travel distance. Additionally, Larp has proven capable of assigning sector-specific risks and generating safety-aware routes at a city-wide scale. The routing solution dynamically adapts maneuverability based on the density of any given region. This success is achieved through the initial decomposition of a potential field into multi-scale cells, with each cell’s restriction zone determined by its proximity to nearby obstacles. To promote transparency and support further research, the source code for an official implementation of Larp has been made publicly accessible at https://github.com/wzjoriv/Larp.
While Larp marks an advancement in UAV route safety, opportunities for improvement remain. The framework’s quad cell decomposition approach and zone classification system are poised for further optimization. The influence of a cell’s dimensions and positioning on the placement of subsequent cells is a critical factor in the pursuit of optimally short and safe routes. Moreover, the current method of zone classification, which relies on a proxy scaled distance to a cell’s periphery to gauge the highest potential across the region, may estimate a significantly higher potential for a region than what it actually is. Future endeavors will concentrate on refining these elements and developing a dynamic and scalable system capable of storing and updating cells in real time to adapt to evolving restrictions for unmanned aircraft system traffic management (UTM).